find the magnitude and direction of the resultant of a 10N-Force that acts angle 40 clockwise from the +x- axis, a 7-N force that acts an angle of 180 clockwise from the + y-axis, 5 -N force that acts at an angle 90 Clockwise from the +y-axis.
1
Expert's answer
2019-09-23T09:09:44-0400

The resultant force


"{\bf R}=\bf F_1+F_2+F_3"

Components of resultanf force


"R_x =F_{1x}+F_{2x}+F_{3x}"

"=10\cos 40^{\circ}+0+5=12.66\:\rm N"

"R_y =F_{1y}+F_{2y}+F_{3y}"

"=-10\sin 40^{\circ}-7+0=-13.43\:\rm N"

The magnitude of a resultant force


"R=\sqrt{R_x^2+R_y^2}=\sqrt{(12.66)^2+(-13.43)^2}=18.5\:\rm N"

The direction of a resultant force


"\theta=\tan^{-1}\left(\frac{R_y}{R_x}\right)"

"=\tan^{-1}\left(\frac{-13.43}{12.66}\right)=46.7^{\circ}\quad ({\rm clockwise\;from\; the}\; +x\:\rm axis)"